(xΒ³-y)dx+xdy=0


1
Expert's answer
2021-10-06T06:16:42-0400

Let us solve the differential equation "(x\u00b3-y)dx+xdy=0," which is equivalent to "xy'-y=-x^3" and to "\frac{y'}x-\frac{y}{x^2}=-x" after dividing by "x^2." It follows that "(\frac{y}x)'=-x," and hence "\frac{y}x=-\frac{x^2}2+C." We conclude that "y=-\frac{x^3}2+Cx" is the general solution of the differential equation.